# Self-Supervised Multi-Part Articulated Objects Modeling via Deformable Gaussian Splatting and Progressive Primitive Segmentation

### v1

@@ Appendix C Experiments Settings
Single-part assumptions (PARIS [20] and ArticulatedGS [21]).
These methods can reconstruct only one movable part at a time. For every movable part, we capture two interaction states in which that part moves while all other parts remain fixed, and we run the pipeline independently for each part. Each state is sampled with 100 random, calibrated RGB views, giving 200 images per part.
Multi-part, two-state methods (DTArt [23] and ArtGS[24]).
These pipelines accept two distinct object states and can handle several movable parts jointly, provided that the part count is known. We therefore collect two interaction states and sample 100 random views per state (total 200 RGB images). The ground-truth number of movable parts is passed to both methods. As DTArt requires depth input, depth maps for every view are rendered in Gazebo [47].
Our method (DeGSS).
DeGSS handles an arbitrary number of parts and interaction states without knowing the part count. To match the total image budget of competing methods, we sample four interaction states in which multiple parts move simultaneously and capture 50 random views per state, producing the same 200-image input. No depth or manual parts information is provided.
@@
Object Method Axis Ang 0\!$\downarrow$ Axis Ang 1\!$\downarrow$ Axis Ang 2\!$\downarrow$ Axis Pos 0\!$\downarrow$ Axis Pos 1\!$\downarrow$ Axis Pos 2\!$\downarrow$ Axis Mot 0\!$\downarrow$ Axis Mot 1\!$\downarrow$ Axis Mot 2\!$\downarrow$
Box 102373 PARIS [20] 9.96 29.61 0.47 0.13 0.31 0.30 30.23 111.21 33.72
ArticulatedGS [21] 0.52 6.53 3.62 0.83 8.32 0.42 5.23 2.13 14.43
DTArt [23] 88.32 62.54 64.68 1.99 5.56 6.79 99.97 67.17 82.00
ArtGS [24] 0.29 0.05 0.21 0.19 0.01 0.39 15.89 0.17 6.02
Ours 0.23 1.34 0.19 0.03 0.23 0.14 1.01 1.59 2.83
Storage 45194 PARIS [20] 64.47 59.08 38.69 0.29 0.13 116.73 109.94 0.32
ArticulatedGS [21] 20.43 4.12 2.32 1.31 0.21 13.11 3.12 1.31
DTArt [23] 27.60 0.20 78.01 4601.75 0.17 59.95 0.16 0.37
ArtGS [24] 54.17 1.68 0.75 0.29 0.08 96.54 1.66 0.02
Ours 9.82 1.56 0.56 0.27 0.03 9.79 1.45 0.07
Storage 45271 PARIS [20] 0.77 87.40 76.58 0.22 0.15 0.17 87.02
ArticulatedGS [21] 1.34 4.24 2.13 0.73 0.42 0.14 0.84
DTArt [23] 4.23 80.96 74.26 12.52 0.01 0.19 50.65
ArtGS [24] 0.02 0.85 0.03 0.14 0.03 0.01 0.02
Ours 0.00 0.34 0.26 0.03 0.00 0.00 0.58
Table 23372 PARIS [20] 1.69 89.40 79.91 0.52 0.84 0.18 157.14 109.55
ArticulatedGS [21] 0.62 5.31 6.24 5.31 0.52 0.54 4.24 13.12
DTArt [23] 0.38 0.22 0.19 0.21 0.01 0.04 0.14 0.48
ArtGS [24] 0.02 89.03 14.64 0.57 1.79 0.02 75.13 13.43
Ours 0.01 0.12 0.29 0.03 0.12 0.00 0.12 0.45
@@
[…]
Box 102373 PARIS [20] 17.70 22.36 279.47 184.97 16.66
ArticulatedGS [21] 16.86 13.59 23.99 5.18 13.23
DTArt [23] 3.88 121.66 199.49 2.25 0.72
ArtGS [24] 9.70 14.18 5.38 2.63 7.78
Ours 6.70 0.88 0.39 2.21 5.75
Storage 45194 PARIS [20] 28.16 8.77 193.38 656.10 176.48
ArticulatedGS [21] 14.48 1.29 0.85 16.15 10.71
DTArt [23] 4.41 71.85 0.09 33.20 1.12
ArtGS [24] 50.88 19.00 1.78 2041.00 6.77
Ours 8.88 0.49 0.51 4.53 6.04
Storage 45271 PARIS [20] 5.09 668.39 703.55 19.35 5.50
ArticulatedGS [21] 14.48 1.28 0.85 16.15 10.71
DTArt [23] 2.94 307.65 F 406.23 8.76
ArtGS [24] 1.40 0.38 0.46 0.14 1.41
Ours 4.24 0.23 0.41 0.23 3.26
Table 23372 PARIS [20] 4.14 5.40 25.67 45.57 2.03
ArticulatedGS [21] 5.72 25.34 0.72 15.96 5.10
DTArt [23] 1.31 0.37 0.18 1.21 0.95
ArtGS [24] 3.38 4.50 1.34 1.05 2.84
Ours 2.17 1.19 0.15 0.25 4.19
Appendix D Additional Results
D.1 Results on Objects with Three Movable Parts
@@
Tab. 2 and Tab. 3 compare all baselines in the three-part instances of PartNet-Mobility [1]. Our method achieves the lowest or second–lowest error in nearly every joint metric (axis angle, axis position, axis motion) and remains highly competitive in the geometry metrics (Chamfer distances). The consistently smaller joint errors indicate that the latent–conditioned deformation field captures the complex motion couplings that arise when three parts move independently, while the geometry scores confirm that the boundary regions between parts are reconstructed sharply.
@@
Axis $\downarrow$ Ang 0 ArtGS [24] 41.86 72.98 54.24 29.24 45.56 63.54 51.24
Ours 0.84 0.32 4.45 1.71 0.88 0.54 1.46
[…]
Ours 0.25 0.34 0.90 1.08 0.58 15.27 3.07
[…]
Ours 0.36 2.07 0.89 1.17 1.12
[…]
Ours 0.32 0.36 0.40 9.31 2.60
[…]
Ours 0.58 0.87 0.63 1.62 9.46 1.79 2.49
[…]
Ours 0.36 0.58 0.74 2.27 4.97 5.04 2.33
CD-s $\downarrow$ ArtGS [24] 41.39 11.38 25.24 53.53 6.34 14.35 25.37
Ours 13.13 6.25 10.19 26.88 4.05 13.36 12.31
CD-m 0 $\downarrow$ ArtGS [24] 121.69 39.09 33.24 104.35 58.24 26.49 63.85
Ours 3.09 4.83 7.19 10.37 5.63 10.08 6.86
CD-m 1 $\downarrow$ ArtGS [24] 180.57 34.18 234.24 156.35 62.64 43.98 118.66
Ours 8.57 3.86 9.41 13.47 0.27 41.44 12.84
CD-w $\downarrow$ ArtGS [24] 7.96 6.09 20.35 96.35 36.25 35.96 33.83
Ours 16.55 3.05 14.80 19.03 12.78 29.90 14.35
D.2 Result on the RS-Art Dataset
@@
Fig. 11 and Tab. 4 present qualitative and quantitative results on our RS-Art dataset. Although the capture setup avoids strong specular highlights and cast shadows, the real objects still exhibit complex materials, weak surface texture, and a narrower viewing envelope than in simulation, making the task substantially harder than on PartNet-Mobility [1]. Among the baselines, only ArtGS can be made to run reliably on the real data, so we compare against it on every instance that yields a converged model. Across all evaluated objects, our method achieves lower joint parameter and geometry errors and produces cleaner part segmentation, confirming that the deformation-field formulation transfers to real scenes. Meanwhile, the results highlight the open challenges discussed in Sec. E: fine surface details degrade under limited texture and grazing angles, and heavy occlusion can still disturb the 3DGS reconstruction. These observations underline the necessity of RS-Art as a bridge between synthetic evaluation and real-world deployment.
Figure 11: Qualitative comparison on real objects from RS-Art.

### v2

@@ Appendix C Experimental Settings
Single-part assumptions (PARIS (Liu et al., 2023) and ArticulatedGS (Guo et al., 2025)). These methods can reconstruct only one movable part at a time. For every movable part, we capture two interaction states in which that part moves while all other parts remain fixed, and we run the pipeline independently for each part. Each state is sampled with 100 random, calibrated RGB views, giving 200 images per part.
Multi-part, two-state methods (DTArt (Weng et al., 2024) and ArtGS(Liu et al., 2025b)). These pipelines accept two distinct object states and can handle several movable parts jointly, provided that the part count is known. We therefore collect two interaction states and sample 100 random views per state (total 200 RGB images). The ground-truth number of movable parts is passed to both methods. As DTArt requires depth input, depth maps for every view are rendered in Gazebo (Koenig and Howard, 2004).
Our method (PD2GS). PD2GS handles an arbitrary number of parts and interaction states without knowing the part count. To match the total image budget of competing methods, we sample four interaction states in which multiple parts move simultaneously and capture 50 random views per state, producing the same 200-image input. No depth or manual parts information is provided.
@@
Appendix D Additional Results
Table 3: Quantitative results with our method restricted to two interaction states as input.
Metric Box Door Eyeglasses Faucet Oven Fridge Storage Table Mean
Axis Ang $\downarrow$ 1.08 8.70 0.35 2.40 1.59 2.64 2.46 3.70 2.87
Axis Pos $\downarrow$ 0.37 1.19 0.26 0.21 0.10 0.24 0.36 0.51 0.41
Part Motion $\downarrow$ 6.86 3.25 7.73 2.96 8.06 9.54 1.74 3.55 5.46
CD-s $\downarrow$ 9.21 6.38 18.55 1.11 2.16 14.31 7.64 4.28 7.96
CD-m $\downarrow$ 6.98 18.06 9.37 0.14 18.78 0.54 13.96 24.85 11.59
CD-w $\downarrow$ 8.90 1.73 11.83 0.54 9.15 10.43 4.67 3.75 6.38
D.1 Results on two interaction states
Because the baselines can use only two states, we limit our method to a total of 200 RGB images as input; our multi-state model therefore receives fewer views per state than the baselines, maintaining a comparable degree of fairness across methods. We add more state = 2 results for direct side-by-side comparison in Tab. 3, further strengthening the fairness of the evaluation. All reported metrics are averaged across all parts and object categories. This experiment extends Tab. 1 setting of main paper to directly assess performance under the two-state constraint. As shown in Tab. 3, our method still maintains strong performance under these settings.
@@ D.1 Results on two interaction states
Object Instance Method Axis Ang 0 $\downarrow$ Axis Ang 1 $\downarrow$ Axis Ang 2 $\downarrow$ Axis Pos 0 $\downarrow$ Axis Pos 1 $\downarrow$ Axis Pos 2 $\downarrow$ Axis Mot 0 $\downarrow$ Axis Mot 1 $\downarrow$ Axis Mot 2 $\downarrow$
Box 102373 PARIS (Liu et al., 2023) 9.96 29.61 0.47 0.13 0.31 0.30 30.23 111.21 33.72
ArticulatedGS (Guo et al., 2025) 0.52 6.53 3.62 0.83 8.32 0.42 5.23 2.13 14.43
DTArt (Weng et al., 2024) 88.32 62.54 64.68 1.99 5.56 6.79 99.97 67.17 82.00
ArtGS (Liu et al., 2025b) 0.29 0.05 0.21 0.19 0.01 0.39 15.89 0.17 6.02
PD2GS (Ours) 0.23 1.34 0.19 0.03 0.23 0.14 1.01 1.59 2.83
Storage 45194 PARIS (Liu et al., 2023) 64.47 59.08 38.69 0.29 0.13 116.73 109.94 0.32
ArticulatedGS (Guo et al., 2025) 20.43 4.12 2.32 1.31 0.21 13.11 3.12 1.31
DTArt (Weng et al., 2024) 27.60 0.20 78.01 4601.75 0.17 59.95 0.16 0.37
ArtGS (Liu et al., 2025b) 54.17 1.68 0.75 0.29 0.08 96.54 1.66 0.02
PD2GS (Ours) 9.82 1.56 0.56 0.27 0.03 9.79 1.45 0.07
Storage 45271 PARIS (Liu et al., 2023) 0.77 87.40 76.58 0.22 0.15 0.17 87.02
ArticulatedGS (Guo et al., 2025) 1.34 4.24 2.13 0.73 0.42 0.14 0.84
DTArt (Weng et al., 2024) 4.23 80.96 74.26 12.52 0.01 0.19 50.65
ArtGS (Liu et al., 2025b) 0.02 0.85 0.03 0.14 0.03 0.01 0.02
PD2GS (Ours) 0.00 0.34 0.26 0.03 0.00 0.00 0.58
Table 23372 PARIS (Liu et al., 2023) 1.69 89.40 79.91 0.52 0.84 0.18 157.14 109.55
ArticulatedGS (Guo et al., 2025) 0.62 5.31 6.24 5.31 0.52 0.54 4.24 13.12
DTArt (Weng et al., 2024) 0.38 0.22 0.19 0.21 0.01 0.04 0.14 0.48
ArtGS (Liu et al., 2025b) 0.02 89.03 14.64 0.57 1.79 0.02 75.13 13.43
PD2GS (Ours) 0.01 0.12 0.29 0.03 0.12 0.00 0.12 0.45
@@
Object Instance Method CD-s $\downarrow$ CD-m 0 $\downarrow$ CD-m 1 $\downarrow$ CD-m 2 $\downarrow$ CD-w $\downarrow$
Box 102373 PARIS (Liu et al., 2023) 17.70 22.36 279.47 184.97 16.66
ArticulatedGS (Guo et al., 2025) 16.86 13.59 23.99 5.18 13.23
DTArt (Weng et al., 2024) 3.88 121.66 199.49 2.25 0.72
ArtGS (Liu et al., 2025b) 9.70 14.18 5.38 2.63 7.78
PD2GS (Ours) 6.70 0.88 0.39 2.21 5.75
Storage 45194 PARIS (Liu et al., 2023) 28.16 8.77 193.38 656.10 176.48
ArticulatedGS (Guo et al., 2025) 14.48 1.29 0.85 16.15 10.71
DTArt (Weng et al., 2024) 4.41 71.85 0.09 33.20 1.12
ArtGS (Liu et al., 2025b) 50.88 19.00 1.78 2041.00 6.77
PD2GS (Ours) 8.88 0.49 0.51 4.53 6.04
Storage 45271 PARIS (Liu et al., 2023) 5.09 668.39 703.55 19.35 5.50
ArticulatedGS (Guo et al., 2025) 14.48 1.28 0.85 16.15 10.71
DTArt (Weng et al., 2024) 2.94 307.65 F 406.23 8.76
ArtGS (Liu et al., 2025b) 1.40 0.38 0.46 0.14 1.41
PD2GS (Ours) 4.24 0.23 0.41 0.23 3.26
Table 23372 PARIS (Liu et al., 2023) 4.14 5.40 25.67 45.57 2.03
ArticulatedGS (Guo et al., 2025) 5.72 25.34 0.72 15.96 5.10
DTArt (Weng et al., 2024) 1.31 0.37 0.18 1.21 0.95
ArtGS (Liu et al., 2025b) 3.38 4.50 1.34 1.05 2.84
PD2GS (Ours) 2.17 1.19 0.15 0.25 4.19
Table 6: Evaluation of our method on object instances with five movable parts. Each reported metric is averaged across individual parts per instance.
Object Instance Method Axis Ang $\downarrow$ Axis Pos $\downarrow$ Part Motion $\downarrow$ CD-s $\downarrow$ CD-m $\downarrow$ CD-w $\downarrow$
Storage 45271 PARIS (Liu et al., 2023) 61.36 0.41 48.79 5.86 380.66 4.96
ArtGS (Liu et al., 2025b) 1.60 0.01 1.94 2.45 4.27 1.77
PD2GS (Ours) 0.02 0.00 0.02 1.39 0.71 1.25
Storage 45612 PARIS (Liu et al., 2023) 61.37 0.14 46.16 5.89 577.17 5.07
ArtGS (Liu et al., 2025b) 0.11 0.08 1.82 2.53 4.09 1.83
PD2GS (Ours) 0.04 0.03 0.01 1.40 0.53 1.32
Table 30666 PARIS (Liu et al., 2023) 36.58 - 0.30 7.73 361.76 6.52
ArtGS (Liu et al., 2025b) 8.91 - 0.15 9.68 1159.34 5.70
PD2GS (Ours) 0.15 - 0.05 4.05 1.47 3.65
Table 33810 PARIS (Liu et al., 2023) 1.07 0.03 0.92 3.53 73.97 3.44
ArtGS (Liu et al., 2025b) 48.67 32.02 8.04 5.75 1346.58 3.18
PD2GS (Ours) 0.44 0.02 0.93 4.85 0.68 3.91
D.2 Results on Objects with Multiple Movable Parts
@@
Tab. 4 and Tab. 5 compare all baselines in the three-part instances of PartNet-Mobility (Xiang et al., 2020). Our method achieves the lowest or second–lowest error in nearly every joint metric (axis angle, axis position, axis motion) and remains highly competitive in the geometry metrics (Chamfer distances). The consistently smaller joint errors indicate that the latent–conditioned deformation field captures the complex motion couplings that arise when three parts move independently, while the geometry scores confirm that the boundary regions between parts are reconstructed sharply.
@@
To further validate scalability to higher articulation complexity, Tab. 6 reports additional results on objects with five movable parts. These instances present substantially more challenging kinematic couplings and occlusion patterns than the two- and three-part cases. Despite the increased difficulty, our approach continues to produce accurate joint estimates and detailed geometry.
Figure 14: Qualitative comparison on real objects from RS-Art.
@@
Axis $\downarrow$ Ang 0 ArtGS (Liu et al., 2025b) 41.86 72.98 54.24 29.24 45.56 63.54 51.24
PD2GS (Ours) 0.84 0.32 4.45 1.71 0.88 0.54 1.46
Axis $\downarrow$ Ang 1 ArtGS (Liu et al., 2025b) 3.43 83.25 23.24 83.24 57.24 24.64 45.84
PD2GS (Ours) 0.25 0.34 0.90 1.08 0.58 15.27 3.07
Axis $\downarrow$ Pos 0 ArtGS (Liu et al., 2025b) 8.89 2.54 352.35 35.77 99.89
PD2GS (Ours) 0.36 2.07 0.89 1.17 1.12
Axis $\downarrow$ Pos 1 ArtGS (Liu et al., 2025b) 8.56 12.23 83.35 45.35 37.37
PD2GS (Ours) 0.32 0.36 0.40 9.31 2.60
Part $\downarrow$ Motion 0 ArtGS (Liu et al., 2025b) 36.75 14.94 9.12 14.56 92.69 14.35 30.40
PD2GS (Ours) 0.58 0.87 0.63 1.62 9.46 1.79 2.49
Part $\downarrow$ Motion 1 ArtGS (Liu et al., 2025b) 16.38 13.07 43.33 63.45 150.35 5.35 48.65
PD2GS (Ours) 0.36 0.58 0.74 2.27 4.97 5.04 2.33
CD-s $\downarrow$ ArtGS (Liu et al., 2025b) 41.39 11.38 25.24 53.53 6.34 14.35 25.37
PD2GS (Ours) 13.13 6.25 10.19 26.88 4.05 13.36 12.31
CD-m 0 $\downarrow$ ArtGS (Liu et al., 2025b) 121.69 39.09 33.24 104.35 58.24 26.49 63.85
PD2GS (Ours) 3.09 4.83 7.19 10.37 5.63 10.08 6.86
CD-m 1 $\downarrow$ ArtGS (Liu et al., 2025b) 180.57 34.18 234.24 156.35 62.64 43.98 118.66
PD2GS (Ours) 8.57 3.86 9.41 13.47 0.27 41.44 12.84
CD-w $\downarrow$ ArtGS (Liu et al., 2025b) 7.96 6.09 20.35 96.35 36.25 35.96 33.83
PD2GS (Ours) 16.55 3.05 14.80 19.03 12.78 29.90 14.35
D.3 Results on the RS-Art Dataset
@@
Fig. 14 and Tab. 7 present qualitative and quantitative results on our RS-Art dataset. Although the capture setup avoids strong specular highlights and cast shadows, the real objects still exhibit complex materials, weak surface texture, and a narrower viewing envelope than in simulation, making the task substantially harder than on PartNet-Mobility (Xiang et al., 2020). Among the baselines, only ArtGS can be made to run reliably on the real data, so we compare against it on every instance that yields a converged model. Across all evaluated objects, our method achieves lower joint parameter and geometry errors and produces cleaner part segmentation, confirming that the deformation-field formulation transfers to real scenes. Meanwhile, the results highlight the open challenges discussed in Sec. E: fine surface details degrade under limited texture and grazing angles, and heavy occlusion can still disturb the 3DGS reconstruction. These observations underline the necessity of RS-Art as a bridge between synthetic evaluation and real-world deployment.
Figure 15: Qualitative comparison on real-world scenes.
Fig. 15 shows qualitative results on RS-Art objects captured in real scenes, where cluttered backgrounds, natural illumination, and real-world materials make reconstruction noticeably more challenging. Even under these conditions, our method recovers sharper geometry and more stable part motions, particularly when object boundaries remain reasonably clear. To ensure a fair evaluation, we also conduct reconstruction comparisons on the real-scene benchmark of the PARIS dataset. In these challenging scenarios where baselines often suffer from severe structural degradation or fail to converge, PD2GS consistently maintains structural integrity and cleaner part boundaries, producing practically usable reconstructions while suppressing artifacts such as surface blurring and deformation spillover.
Table 8: Robustness evaluation under camera pose inaccuracies.
Method Axis Ang $\downarrow$ Axis Pos $\downarrow$ Part Motion $\downarrow$ CD-s $\downarrow$ CD-m $\downarrow$ CD-w $\downarrow$
PARIS (Liu et al., 2023) w/o Perturbation 29.98 0.38 70.84 12.98 103.22 10.95
ArtGS (Liu et al., 2025b) w/o Perturbation 3.53 3.46 5.49 4.29 82.65 6.72
PD2GS (Ours) w/ (1.0°, 0.02) 2.92 0.22 3.66 11.02 11.09 7.05
PD2GS (Ours) w/ (3.0°, 0.05) 10.00 0.24 1.55 13.60 24.71 9.19
PD2GS (Ours) w/ colmap 1.93 0.11 3.85 7.39 6.82 7.63
PD2GS (Ours) w/o Perturbation 0.40 0.06 0.43 2.05 1.12 3.02
D.4 Effect of Camera Pose Errors
To evaluate robustness to camera pose inaccuracies, we design three controlled perturbation settings. Starting from ground-truth poses, we inject Gaussian noise into the rotation and translation components to form two perturbation levels, denoted as $(u,v)$, where $u$ is the standard deviation of rotational noise measured in degrees and $v$ is that of translational noise measured in scene units. In addition, we test a third setup that uses camera poses estimated by COLMAP. All methods are evaluated using the same protocol as Tab. 1 in the main paper. Metrics are averaged across all articulated parts and objects. As shown in Tab. 8, although performance decreases slightly due to these perturbations, our method consistently maintains superior results compared to prior methods.
Table 9: Quantitative comparison with articulated generation baselines on the Storage category.
Object Method Axis Ang $\downarrow$ Axis Pos $\downarrow$ Part Motion $\downarrow$ CD-s $\downarrow$ CD-m $\downarrow$ CD-w $\downarrow$
Storage NAP (Lei et al., 2023) 3.88 2.45 7.29 5.56 5.19 8.54
SINGAPO (Liu et al., 2025a) 2.72 0.86 9.03 - - -
PD2GS (Ours) 0.13 0.02 0.32 2.78 0.07 5.48
D.5 Comparison with articulated generation methods
We further extended our evaluation to include generative articulated object models, NAP (Lei et al., 2023) and SINGAPO (Liu et al., 2025a). NAP (Lei et al., 2023) operates by generating target objects from Gaussian noise derived from specific instances, whereas SINGAPO (Liu et al., 2025a) infers them from a single input view. To ensure a robust evaluation for SINGAPO (Liu et al., 2025a) given its single-view dependence, we randomly sampled 50 input views for each object and reported the metrics from the best-performing result. As shown in Tab. 9, both methods exhibit significant deviations in joint motion analysis compared to our approach, reflecting the inherent distinction between generative plausibility and the precise kinematic inference required for our task.
Table 10: Training and inference time for objects with varying numbers of articulated parts.
Number of parts Method 2 3 5
Training (min) PARIS (Liu et al., 2023) 31.38 45.99 75.15
DTArt (Weng et al., 2024) 26.31 31.42 43.76
ArtGS (Liu et al., 2025b) 16.22 15.08 17.86
PD2GS (Ours) 35.59 38.20 44.82
Inference (ms) PARIS (Liu et al., 2023) 24773.60 37180.59 61924.30
DTArt (Weng et al., 2024) 568.67 525.70 613.30
ArtGS (Liu et al., 2025b) 3.42 3.71 4.16
PD2GS (Ours) 37.93 40.86 37.22
D.6 Results on Training and Inference Time
We provide statistics on training and inference overhead for reference, as shown in Tab. 10. For a fair comparison, all methods are trained on the same dataset of 200 views per object. Regarding inference, since discrete-state baselines (DTArt (Weng et al., 2024), ArtGS (Liu et al., 2025b)) cannot natively infer arbitrary continuous states, we measured their timings by explicitly rigidly transforming their reconstructed models using inferred joint parameters. Specifically, for ArtGS (Liu et al., 2025b), we render by rigidly transforming the reconstructed Gaussian field; thus, the inference time is equivalent to standard static 3DGS (Kerbl et al., 2023). In contrast, our method requires the evaluation of a latent-conditioned deformation field. In the context of articulated object modeling, inference efficiency is not the primary limiting factor. Thus, this difference is secondary to the substantial gains in modeling capability.
D.7 Ablation Studies